Logiweb dictionary of append

Up Help

0 0 append

1 2 append(∗, ∗)

The pyk compiler, version 0.grue.20050502+ by Klaus Grue GRD-2005-05-26.UTC:20:37:04.090427 = MJD-53516.TAI:20:37:36.090427 = LGT-4623856656090427e-6